a number is 7 less than another, if 4 times the smaaler number plus twice the larger is 46. find the two numbers?: a number is 7 less than another, if 4 times the smaaler number plus twice the larger is 46. find the two numbers?
Answer by ptaylor(1256)


Let x=larger number
Then 2x=twice the larger number
And let x-7=smaller number
Then 4(x-7)=four times the smaller number
Now we are told that;
4(x-7)+2x=46 get rid of parens (distributive law)
4x-28+2x=46 add 28 to both sides
4x-28+28+2x=46+28 collect like terms
6x=74
x= 12 1/3-------------larger number
x-7=(12 1/3)-7=5 1/3----------------smaller number
CK
2*(12 1/3)+4*(5 1/3)=46
2*(37/3)+4(16/3)=46
74/3+64/3=46
138/3=46
46=46
Hope this helps---ptaylor